You are dealt two playing cards from a shuffled​ 52-card deck. Per​ usual, this deck contains 4 aces. What is the probability that both of your cards are​ aces?
Answer #1

2 cards from 52 cards can be selected in 52C2 ways

2 aces from 4 can be selected in 4C2 ways

Required probability = 4C2/52C2 = 6/1326 = 0.004525